OpenStudy (anonymous):

can you please help me solve this quadratic equation for x by factoring- 2x^2+(2x+4)^2=(2x+6)^2

OpenStudy (anonymous):

x^2+(2x+4)^2=(2x+6)^2

OpenStudy (rajee_sam):

first do FOIL for the (2x + 2)^2 and (2x + 6)^2

OpenStudy (rajee_sam):

Then combine like terms and form one single quadratic equation and then factor

OpenStudy (rajee_sam):

\[x ^{2} + (2x + 4)^{2} = (2x + 6)^{2}\] \[x^{2} + (2x + 4) (2x + 4) = (2x + 6) (2x + 6)\] \[x ^{2} + 4x ^{2} + 8x + 8x + 16 = 4x ^{2} + 12x + 12x + 36\] \[x ^{2} = 4x + 4x + 20\] \[x ^{2} - 8x - 20 = 0\] \[(x - 10) (x + 2) = 0\] \[x = 10 ; x = -2\]